OpenStudy (anonymous):

Please help me! Will fan and medal! 1)In two or more complete sentences, explain how to find the exact value of secant(13pi/6) including quadrant location. 2)In two or more complete sentences, explain how to find the exact value of cotangent (7pi/4) including quadrant location.

OpenStudy (welshfella):

13pi/6 - 2pi = 13pi/6 - 12pi/6 = pi/6 which is an angle in the first quadrant. You can find the cosine of pi/6 from the standard 30/60/90 triangle . The secant is = 1 / cosine.

OpenStudy (welshfella):